Jennifer needs to make a solution that is 30% hydrochloric acid. If she must use all of the 25 mL of 60% hydrochloric acid she started with, then how much water needs to be mixed with the original solution?

Answer by Alan3354(69443)

Since 30% is 1/2 of 60%, just add an equal amount of water.
25 mL of water.
-------------
Or,
0.6*25 + 0*x = 0.3*(25+x)
15 = 7.5 + 0.3x
0.3x = 7.5
x = 25 mL